OpenStudy (anonymous):

In a rush, help anyone? The number of telephones in a certain town is 48,000. If this represent 12.8 telephones per 100 of population, the population of this town to the nearest thousand is.. (A) 128,000 (B) 375,000 (C) 378,000 (D) 566,000 (E) 560,000 Is it A?

Directrix (directrix):

I am tempted to look for meaning in this proportion: 12.8 is to 100 as P is the 48000 where P is the population of the town. Actually, we don't know that everyone in town has a phone so the value of P is, I think, the number of people in the town who have phones. 12.8/100 = P/48000 Cross multiply and solve for P and see what pops up. @KawasumiKimiko

OpenStudy (anonymous):

I got p=6000

Directrix (directrix):

I got 6144 which cannot be correct. I think the proportion I posted is messed up. 12.8 is to 100 as 48000 is to P where P is the population of the town. 12.8/100 = 48000/P See what P is.

Directrix (directrix):

I think we have a winner. What did you get?

OpenStudy (anonymous):

I got 384,000

Directrix (directrix):

I wonder why I got 375,000. http://www.wolframalpha.com/input/?i=12.8%2F100+%3D+48000%2Fx

OpenStudy (anonymous):

Oh..I think i did a mistake in my dividing..